OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (2, 7) and (-2, 9).

OpenStudy (anonymous):

y=1/2-8 y=1/2+8 y=-1/2-8 y=-1/2+8

OpenStudy (anonymous):

first draw a graph

OpenStudy (inkyvoyd):

Find the slope of the line.

OpenStudy (anonymous):

is it c

OpenStudy (inkyvoyd):

No. Where'd you get c from?

OpenStudy (anonymous):

like option c?

OpenStudy (inkyvoyd):

But why C?

OpenStudy (inkyvoyd):

Why not A? or B? or D?

OpenStudy (anonymous):

becasue th ats th eanswer i got ?

OpenStudy (inkyvoyd):

Show work?

OpenStudy (anonymous):

y= -1/2x+8

OpenStudy (inkyvoyd):

@sammy12 , that's incorrect. The answer is C

OpenStudy (anonymous):

so im right

OpenStudy (inkyvoyd):

Actually, you don't know who's right.

OpenStudy (inkyvoyd):

Wanna listen to my explanation?

OpenStudy (inkyvoyd):

@needhomeworkhelp

OpenStudy (anonymous):

yeeess!!!!

OpenStudy (inkyvoyd):

Ok. Slope=rise/run

OpenStudy (inkyvoyd):

Since we have two points we search for the slope first.

OpenStudy (inkyvoyd):

(2, 7) and (-2, 9). we have rise (9-7) rise=2 run (-2-2) run=-4 slope=rise/run=2/(-4)=\(-\frac{1}{2}\) So, we know the slope is 1/2. What now?

OpenStudy (inkyvoyd):

We have a linear equation y=mx+b slope is m We don't know b. But, we know a 2 points, and the slope of the line. Just pick a point (x,y) and put it in there, as well as the slope m, to find b.

OpenStudy (inkyvoyd):

\(\large y=-\frac{1}{2}x+b\) Let's use the point (2,7) 7=-1+b b=8 So, the answer was actually \(\Huge B.\)

OpenStudy (inkyvoyd):

*the answer was \(\Huge D\), not \(\large B\).